OpenStudy (maddytarin_98):

An experienced computer technician can set up a network in 12 hours. It would take an intern 20 hours to set up the same network. If they work together, how long will it take to set up the network? The rate of the expierenced technician is 1/12 It takes ______ hours for the experienced technician and the intern to set up the network if they work together.

OpenStudy (kl0723):

for this problem you are given a hint and that's the rate at which the experienced worker does the job = 1/12

OpenStudy (kl0723):

what would be the rate at which the intern does the job?

OpenStudy (maddytarin_98):

that is unknown ?

OpenStudy (kl0723):

if the experienced person can do 1 network in 12 hours his rate is 1/12

OpenStudy (kl0723):

so the intern can do 1 network setup in 20 hours, his rate will be 1/20

OpenStudy (kl0723):

you can add both rates 1/12+1/20 = 1/t

OpenStudy (maddytarin_98):

ohh okay so thats what i should fill in?

OpenStudy (kl0723):

why 1/t? because they complete 1 network in (t) amount of time

OpenStudy (kl0723):

did you add the rates? that should be equal to 1/t

OpenStudy (maddytarin_98):

no i didn't !

OpenStudy (maddytarin_98):

so t=1/32 ?

OpenStudy (wolf1728):

No Add 1/12 plus 1/20

OpenStudy (kl0723):

you should get that 1/12 + 1/20 = 2/15

OpenStudy (kl0723):

and that addition of rates is equal to 1 network setup in a given time (t)

OpenStudy (kl0723):

2/15 = 1/t

OpenStudy (kl0723):

makes sense?

OpenStudy (maddytarin_98):

im so confused on whats the answer ;(

OpenStudy (kl0723):

the answer is t we're trying to solve for it

OpenStudy (kl0723):

was the addition of the rates confusing?

OpenStudy (maddytarin_98):

yeah I'm on a timed quizzes and I appreciate the help and interactiveness, but I'm looking for answers :( thank you so much id understand if you didn't want to help anymore

OpenStudy (kl0723):

I understand, so we had that 2/15 = 1/t

OpenStudy (kl0723):

you can now flip both sides of the equation and isolate t

OpenStudy (kl0723):

\[t = \frac{ 15 }{ 2}\]

OpenStudy (kl0723):

what's 15/2 in hours?

OpenStudy (wolf1728):

In one hour 2/15 of a network is set up. 2/15 Network/Hour We need to know how many hours per network So it becomes 15/2 hours

OpenStudy (kl0723):

exacly so t = 15/2 = 7.5 hours

OpenStudy (maddytarin_98):

THANK YOU ALL FOR HELPING me 7.5 was correct !!!

OpenStudy (wolf1728):

Thanks maddy and good work kI0723

OpenStudy (kl0723):

np, please take some time to understand what's going on, it's always important for future courses

OpenStudy (kl0723):

thanks @wolf1728

OpenStudy (wolf1728):

thanks - you did very well

OpenStudy (maddytarin_98):

yeah! I totally agree! its just that my summer school is due tomorrow and i have so much to do and i know i sound greedy but i really need answers :(